# Benchmark for Antibody Binding Affinity Maturation and Design

### v1

@@ Abstract
We introduce AbBiBench (Antibody Binding Benchmarking), a benchmarking framework for antibody binding affinity maturation and design. Unlike existing antibody evaluation strategies that rely on antibody alone and its similarity to natural ones (e.g., amino acid recovery rate, structural RMSD), AbBiBench considers an antibody–antigen (Ab-Ag) complex as a functional unit and evaluates the potential of an antibody design binding to given antigen by measuring protein model’s likelihood on the Ab-Ag complex. We first curate, standardize, and share 9 datasets containing 9 antigens (involving influenza, anti-lysozyme, HER2, VEGF, integrin, and SARS-CoV-2) and 155,853 heavy chain mutated antibodies. Using these datasets, we systematically compare 14 protein models including masked language models, autoregressive language models, inverse folding models, diffusion-based generative models, and geometric graph models. The correlation between model likelihood and experimental affinity values is used to evaluate model performance. Additionally, in a case study to increase binding affinity of antibody F045-092 to antigen influenza H1N1, we evaluate the generative power of the top-performing models by sampling a set of new antibodies binding to the antigen and ranking them based on structural integrity and biophysical properties of the Ab-Ag complex. As a result, structure-conditioned inverse folding models outperform others in both affinity correlation and generation tasks. Overall, AbBiBench provides a unified, biologically grounded evaluation framework to facilitate the development of more effective, function-aware antibody design models.
@@ 1 Introduction
Antibodies are critical components of the adaptive immune system, functioning primarily by recognizing and binding specifically to antigens such as pathogens or aberrant cells. This specific recognition is facilitated through complementary regions: the antibody’s paratope and the antigen’s epitope. Enhancing the binding affinity between an antibody and its corresponding antigen can significantly increase therapeutic potency and diagnostic sensitivity, making antibody design and optimization a significant area of drug discovery. Developing a therapeutic monoclonal antibody depends on multiple factors, including expression, stability, immunogenicity, aggregation, and binding affinity [9, 30]. Among these, binding affinity is the most critical determinant of therapeutic efficacy, as it directly influences antibody potency. Insufficient binding affinity fundamentally limits an antibody’s utility. Therefore, increasing binding affinity between antibody and antigen has been a crucial process in therapeutic antibody development.
Traditional antibody discovery methods, such as phage display technology [48, 50, 65] and animal immunization [21, 37], employ iterative cycles of mutation and selection to progressively improve binding affinity. These methods have significantly advanced therapeutic antibody discovery but face challenges due to the initially limited diversity of libraries in phage display or the narrow naive B cell repertoire available in animal immunization models, restricting the comprehensive exploration of potential high-affinity antibody variants (Fig. 1). Machine learning-based antibody design approaches complement these experimental techniques by efficiently navigating the vast search space (Fig. 1) and proposing high-affinity antibody variants that may not be readily accessible through experimental approaches alone.
@@
Recent advances in machine learning, especially with protein language models and structure-based generative models [12, 24, 26, 27, 38, 39, 41, 45, 47, 49, 61, 63, 67, 75, 78], have shown promise in antibody design. However, common evaluation metrics like amino acid recovery rates or structural RMSD to natural antibodies do not adequately capture biological relevance in antibodies. In general protein design, comparing to the closest natural variant is a reasonable validation strategy because protein mutations are driven by strong evolutionary pressure. In contrast, antibodies are uniquely generated through semi-random process such as V(D)J recombination and somatic hypermutation, resulting in extraordinarily high sequence diversity. It is common for two antibodies that bind the same antigen to share minimal sequence similarity, and exceedingly rare that two arbitrary antibodies are similar unless they are from the same lineage of a human immune response. As such, evaluating designed antibodies by how closely they resemble naturally occurring ones overlooks the fundamental biology of antibody generation. This calls for new evaluation criteria that better reflect the functional goals of antibody design, rather than assumptions borrowed from general protein modeling.
@@
From a structural biology perspective, binding affinity is determined not just by the antibody sequence, but by the quality of the interface it forms with the antigen. High-affinity binding typically arises from antibody-antigen (Ab-Ag) complexes that exhibit structural integrity [63] – meaning they are stable, well-packed, and maintain favorable conformations with minimal strain. Structural integrity ensures optimal shape and chemical complementarity at the binding interface. Antibodies that form such stable complexes with their targets resemble naturally occurring Ab-Ag complexes collected in structural database SAbDab [15]. Recent machine learning models [12, 27, 41, 45, 47, 67, 78] have shown success in learning Ab-Ag complex sequence-structure patterns, enabling us to gauge whether a designed Ab-Ag complex lies within the high-probability manifold of structurally stable, high-affinity complexes. Therefore, incorporating the antigen into evaluation provides a more biologically grounded and functionally relevant assessment of antibody design.
@@
To address above limitations, we introduce AbBiBench (Antibody Binding Benchmarking), a biologically relevant benchmarking framework specifically designed for antibody binding affinity maturation. Rather than assessing antibodies in isolation [9], we consider the Ab-Ag complex as a unit for evaluation. We curated standardized data from publicly available experimental binding affinity studies, compiling 155,853 mutated heavy chain antibodies across nine antigen targets to evaluate protein models for binding affinity optimization. [^1] https://huggingface.co/datasets/AbBibench/Antibody_Binding_Benchmark_Dataset We also devised and publicly shared an efficient pipeline to rank newly designed antibodies based on complex structural integrity and biophysics properties. [^2] https://github.com/MSBMI-SAFE/AbBiBench AbBiBench is curated to avoid data leakage: although wild-type antibodies or antigens may appear in public datasets, no training corpus contains the mutant antibody–antigen complexes it evaluates. By providing a rigorous, biologically grounded benchmark for antibody design, AbBiBench will accelerate methods that lead to clinically and diagnostically impactful discoveries.
@@ 2.1 Measuring Binding Affinity Between Antibody and Antigen
Three main approaches exist: experimental [1, 33, 43], biophysics-based [6, 8, 76], and data-driven (machine learning) methods [12, 18, 34, 42]. The most direct method to measure binding affinity is through the dissociation constant (Kd), which experimentally quantifies how tightly an antibody binds to its antigen. They are costly and low-throughput, thus computational models aim to approximate the binding affinity to enable large-scale screening. Biophysics-based computational models estimate Ab-Ag binding affinity by calculating interaction energies and free energy change, such as FoldX and $\Delta\Delta G$. It provides mechanistic insights into binding affinity. Though not direct affinity estimator, structural integrity (or plausibility) of Ab-Ag complex have been used as a proxy for binding affinity, measured via structure-based confidence metrics (interface pLDDT, iPTM) [18, 34] or ML-based perplexity (e.g., perplexity from AntiBERTy, ProteinMPNN, or ESM-IF) [12, 42, 61]. Despite the efforts to measure binding affinity computationally, the reliability and accuracy of these metrics remain an ongoing challenge [32].
@@ 2.2 Prior Benchmark Studies Evaluating Binding Affinity via Model Likelihood
Alternatively, some studies retrospectively evaluate machine learning models using historical experimental protein fitness data [9, 70, 72]. These works assume that higher model likelihoods correlate with higher protein fitness, suggesting the model can generate functional sequences. ProteinGym [55] exemplifies this approach by benchmarking models against experimental data on enzymatic activity, binding, expression, and stability. It measures correlation between model perplexity and fitness scores. However, ProteinGym does not include antibodies. FLAb [9] addresses this gap by compiling experimental measurements specifically for antibodies, including binding affinity, specificity (e.g., polyreactivity), immunogenicity, and developability metrics. While comprehensive, FLAb does not consider the antigen when evaluating binding affinity. This is a critical limitation, as antibody-antigen interactions are highly specific, and accurate affinity prediction requires modeling the Ab-Ag complex.
@@ 3.1 Datasets
We compiled 11 datasets of antigen, antibody heavy chain sequences, antibody structure, and experimental binding affinity measurements (Table 1). We focus on the mutation in variable heavy chain (vH) due to its high diversity from V(D)J recombination and the central role of CDR-H3 in antigen binding [7]. The vH often acts as a unique antigen-specific signature [13], while the light chain remains relatively conserved across functional antibodies [29], making vH the most relevant region for affinity optimization. For antibody structure, we focus on variable regions of heavy and light chain, where affinity conferring mutations occurs. We only included datasets with at least 20 mutated antibodies to ensure the statistical significance of evaluation.
In each binding affinity study, antibody libraries are constructed through phage or yeast display, introducing mutations via deep mutational scanning or at targeted positions. Some libraries were designed computationally using machine learning [40, 62] or biophysical modeling [10], resulting in variant sets ranging from 40 to 65,535. For consistency, we standardized experimental affinity measurements (Kd and enrichment ratio) by taking the negative log (for Kd) or log (for enrichment), so that higher values indicate stronger binding (Supplement LABEL:supplementary:dataset, Table LABEL:tab:dataset_details). While log enrichment is an indirect measure, log enrichment reflects how well a variant is retained after antigen-specific selection and correlates with binding strength. When normalized, it provides a scalable proxy for relative binding affinity in high-throughput screens. Notably, AbBiBench is designed to avoid data leakage: while wild-type antibodies and antigens may occur in public datasets such as OAS [57] and SAbDab [15], the specific antibody mutant–antigen complexes curated in our benchmark are not present in any known training corpus.
PDB ID Antibody Antigen Heavy-chain variants Binding score Study 4fqi_h1 CR9114 Influenza A/ New Caledonia/20/99 (H1N1) $65,094$ $-\!\log K_{d}$ [60] 4fqi_h3 CR9114 Influenza A/ Wisconsin/67/2005 (H3N2) $65,535$ $-\!\log K_{d}$ [60] 3gbn_h1 CR6261 Influenza A/ New Caledonia/20/99 (H1N1) $1,887$ $-\!\log K_{d}$ [60] 3gbn_h9 CR6261 Influenza A/ Hong Kong/1073/1999 (H9N2) $1,842$ $-\!\log K_{d}$ [60] aayl49 AAYL49 Spike HR2 $4,312$ $-\!\log K_{d}$ [17] aayl49_ML AAYL49_ML Spike HR2 $8,953$ $-\!\log K_{d}$ [40] aayl51 AAYL51 Spike HR2 $4,320$ $-\!\log K_{d}$ [17] 2fjg G6.31 VEGF $2,223$ log enrichment [36] 1mlc D44.1 Hen-egg-white lysozyme $1,297$ log enrichment [74] 1n8z trastuzumab HER2 $419$ $-\!\log K_{d}$ [62] 1mhp AQC2 Integrin-$\alpha$-1 $40$ $-\!\log K_{d}$ [10]
Table 1: Overview of the 11 Ab-Ag binding-affinity assays reported in AbBiBench, showing the number of heavy-chain mutants in each study and the respective binding metric.
@@ 3.2 Protein models
Masked Language Models: Masked protein language models (MLMs) predict masked residues based on context and capture correlations between sequence motifs and higher-level functional properties, enabling their application in antibody design [25, 51]. We evaluate several representative protein MLMs. ESM-2 [42] is trained on large-scale protein sequence datasets using a masked language modeling objective. AntiBERTy [61] is a 12-layer BERT model trained on 57 million heavy- and light-chain sequences from antibody database (OAS [57]). Incorporating structural information into protein language models (PLMs) improves their ability to capture spatial context beyond sequence proximity. SaProt [67] extends ESM-2 with structure-aware tokens from Foldseek [4, 71], embedding residue identity and local structure. ProSST [41] uses geometric vector perceptrons (GVP) encoder [31] that discretizes local atomic neighborhoods into a compact codebook, with disentangled attention over sequence, structure, and position. ESM-3 [24], an upgraded ESM-2, is a multimodal PLM that models sequence, structure, and function through discrete token tracks. Based on such general-purpose protein MLMs, several antibody-specific MLMs have been developed, including CurrAb [5], AbLang [58], and IgBlend [47], to capture antibody-specific mutations driven by somatic recombination and hypermutation [61]. CurrAb, a fine-tuned ESM-2, uses curriculum learning to gradually shift from unpaired to paired OAS [57] antibody data while preserving pre-trained knowledge on general proteins.
@@
Autoregressive Protein Language Models: Unlike MLMs, which predict masked tokens based on bidirectional context, autoregressive PLMs generate the next token using only left-to-right context. ProGen2 [54] is a Transformer-decoder model that scales up to 6.4 B parameters and is trained on $\sim\!1$ billion natural protein sequences. ProtGPT2 [19] adopts the GPT-2 architecture with 738 M parameters and is trained end-to-end on $\sim\!50$ million protein sequences. Inverse folding models are also autoregressive but based on global structure embeddings. Therefore, ProGen2 and ProtGPT2 can serve as structure-agnostic comparisons for inverse folding models.
@@
Inverse Folding Models: Inverse folding models aim to predict amino acid sequences from a given protein structure, often in an autoregressive manner. This approach has been applied to antibody mutant design by leveraging known Ab-Ag complex structures [63], enabling sequence exploration to identify mutations that preserve or enhance complex stability and binding affinity. Widely used models include ProteinMPNN [12] and ESM-IF [27]. ProteinMPNN uses a message-passing neural network [20] to model residue interactions before autoregressive sequence generation. ESM-IF1 combines a GVP encoder for extracting backbone-invariant features with a Transformer decoder. AntiFold [26], based on ESM-IF1, is fine-tuned on antibody structures from SAbDab and OAS.
@@
Diffusion-Based Generative Models: Diffusion models approach antibody design as a denoising process, transforming Gaussian noise into a target antibody by learning the joint distribution of atomic coordinates, orientations, and residue identities. DiffAb [45] conditions on an antigen–antibody framework complex and jointly diffuses CDR sequence and structure. To assess backbone flexibility, we also evaluated fixed-backbone variants DiffAb_fixbb, a common setting in protein design [3, 27, 28, 45, 66, 69]. AbDiffuser [49] extends to full-atom generation with physical priors for side chains. IgDiff [11] performs de novo backbone generation by sampling variable-region backbones and then fills in the sequences using AbMPNN [14].
@@
CDR Imputation in Geometric Representation: Geometry-aware methods view affinity maturation as filling in missing CDRs within the explicit 3-D Ab-Ag interface. MEAN [38] masks CDRs on an E(3)-equivariant residue–atom graph containing both chains and epitope; two alternating message-passing blocks jointly restore CDR sequence and backbone. dyMEAN [39] upgrades this to full-atom, end-to-end design: conserved-framework initialization plus a “shadow paratope” lets the network emit paratope sequence, side-chain geometry, and binding pose in one shot. We also considered fixed-backbone versions of MEAN and dyMEAN, dubbed MEAN_fixbb and dyMEAN_fixbb, respectively.

### v2

# AbBiBench: A Benchmark for Antibody Binding Affinity Maturation and Design

@@ Abstract
We introduce AbBiBench (Antibody Binding Benchmarking), a benchmarking framework for antibody binding affinity maturation and design. Unlike previous strategies that evaluate antibodies in isolation, typically by comparing them to natural sequences with metrics such as amino acid recovery rate or structural RMSD, AbBiBench instead treats the antibody–antigen (Ab–Ag) complex as the fundamental unit. It evaluates an antibody design’s binding potential by measuring how well a protein model scores the full Ab–Ag complex. We first curate, standardize, and share more than 184,500 experimental measurements of antibody mutants across 14 antibodies and 9 antigens—including influenza, lysozyme, HER2, VEGF, integrin, Ang2, and SARS-CoV-2—covering both heavy-chain and light-chain mutations. Using these datasets, we systematically compare 15 protein models including masked language models, autoregressive language models, inverse folding models, diffusion-based generative models, and geometric graph models by comparing the correlation between model likelihood and experimental affinity values. Additionally, to demonstrate AbBiBench’s generative utility, we apply it to antibody F045-092 in order to introduce binding to influenza H1N1. We sample new antibody variants with the top-performing models, rank them by the structural integrity and biophysical properties of the Ab–Ag complex, and assess them with in vitro ELISA binding assays. Our findings show that structure-conditioned inverse folding models outperform others in both affinity correlation and generation tasks. Overall, AbBiBench provides a unified, biologically grounded evaluation framework to facilitate the development of more effective, function-aware antibody design models.
@@ 1 Introduction
Antibodies are critical components of the adaptive immune system, functioning primarily by recognizing and binding specifically to antigens such as pathogens or aberrant cells. This specific recognition is facilitated through complementary regions: the antibody’s paratope and the antigen’s epitope. Improving antibody–antigen affinity boosts therapeutic potency and is critical for drug discovery. Developing a therapeutic monoclonal antibody depends on multiple factors, including expression, stability, immunogenicity, aggregation, and binding affinity Chungyoun et al. (2024); Jain et al. (2017). Among these, binding affinity is the most critical determinant of therapeutic efficacy, as it directly influences antibody potency. Therefore, increasing binding affinity between antibody and antigen has been a crucial process in therapeutic antibody development.
Traditional antibody discovery methods, such as phage display technology Marks et al. (1992); McCafferty et al. (1990); Smith (1985) and animal immunization Green et al. (1994); Köhler & Milstein (1975), employ iterative cycles of mutation and selection to progressively improve binding affinity. These methods have significantly advanced therapeutic antibody discovery but face challenges due to the initially limited diversity of libraries in phage display or the narrow naive B cell repertoire available in animal immunization models, restricting the comprehensive exploration of potential high-affinity antibody variants (Fig. 1). Machine learning-based antibody design approaches complement these experimental techniques by efficiently navigating the vast search space (Fig. 1) and proposing high-affinity antibody variants that may not be readily accessible through experimental approaches alone.
@@
Recent advances in machine learning, especially with protein language models and structure-based generative models Dauparas et al. (2022); Hayes et al. (2025); Høie et al. (2025); Hsu et al. (2022); Kong et al. (2022; 2023); Li et al. (2024); Luo et al. (2022); Malherbe & Ucar (2024); Martinkus et al. (2023); Ruffolo et al. (2021); Shanker et al. (2024); Su et al. (2023); Watson et al. (2023); Wu & Li (2023), have shown promise in antibody design. However, common evaluation metrics like amino acid recovery rates or structural RMSD to natural antibodies do not adequately capture biological relevance in antibodies. In general protein design, comparing to the closest natural variant is a reasonable validation strategy because protein mutations are driven by strong evolutionary pressure. By contrast, antibody generation involves stochastic recombination and hypermutation, yielding extreme diversity. Even antibodies to the same antigen often show little sequence similarity unless clonally related. As such, evaluating designed antibodies by how closely they resemble naturally occurring ones overlooks the fundamental biology of antibody generation. This calls for new evaluation criteria that better reflect the functional goals of antibody engineering, rather than assumptions borrowed from general protein design.
@@
From a structural biology perspective, binding affinity is determined not just by the antibody sequence, but by the quality of the interface it forms with the antigen. High-affinity binding typically arises from antibody-antigen (Ab-Ag) complexes that exhibit structural integrity Shanker et al. (2024) – meaning they are stable, well-packed, and maintain favorable conformations with minimal strain. Structural integrity ensures optimal shape and chemical complementarity at the binding interface. Antibodies that form such stable complexes with their targets resemble naturally occurring Ab-Ag complexes such as those collected in structural databases like SAbDab Dunbar et al. (2014). Recent machine learning models Dauparas et al. (2022); Hsu et al. (2022); Li et al. (2024); Luo et al. (2022); Malherbe & Ucar (2024); Su et al. (2023); Wu & Li (2023) have shown success in learning Ab‑Ag complex sequence-structure patterns, enabling us to gauge whether a designed Ab-Ag complex lies within the high‑probability manifold of structurally stable, high‑affinity complexes. Therefore, incorporating the antigen into evaluation provides a more biologically grounded and functionally relevant assessment of antibody design.
@@
To address the limitations discussed above, we introduce AbBiBench (Antibody Binding Benchmarking), a biologically relevant benchmarking framework specifically designed for improving antibody binding affinity. Rather than assessing antibodies in isolation Chungyoun et al. (2024), we consider the Ab-Ag complex as a unit for evaluation. We curated standardized data from publicly available experimental binding affinity studies, compiling 184,500 mutated antibodies across nine antigen targets to evaluate protein models for binding affinity optimization. [^1] https://huggingface.co/datasets/AbBibench/Antibody_Binding_Benchmark_Dataset We also devised and publicly shared an efficient pipeline to rank newly designed antibodies based on complex structural integrity and biophysical properties. [^2] https://github.com/MSBMI-SAFE/AbBiBench AbBiBench is curated to avoid data leakage: although wild-type antibodies or antigens may appear in public datasets, no training corpus contains the mutant antibody–antigen complexes it evaluates. By providing a rigorous, biologically grounded benchmark for antibody design, AbBiBench will accelerate methods that lead to clinically and diagnostically impactful discoveries.
@@ 2.1 Measuring Binding Affinity Between Antibody and Antigen
Three main approaches for measuring binding affinity exist: experimental Abdiche et al. (2008); Jönsson et al. (1991); Livingstone (1996), biophysics-based Buß et al. (2018); Chi et al. (2024); Weitzner et al. (2017), and data-driven (machine learning) methods Dauparas et al. (2022); Evans et al. (2021); Jumper et al. (2021); Lin et al. (2023). The most direct method to measure binding affinity is through the dissociation constant (Kd), which experimentally quantifies how tightly an antibody binds to its antigen. They are costly and low-throughput, thus computational models aim to approximate the binding affinity to enable large-scale screening. Biophysics-based computational models estimate Ab-Ag binding affinity by calculating interaction energies and free energy change, such as FoldX and $\Delta\Delta G$. It provides mechanistic insights into binding affinity. Though not direct affinity estimator, structural integrity (or plausibility) of Ab-Ag complex have been used as a proxy for binding affinity, measured via structure-based confidence metrics (interface pLDDT, iPTM) Evans et al. (2021); Jumper et al. (2021) or ML-based perplexity (e.g., perplexity from AntiBERTy, ProteinMPNN, or ESM-IF) Dauparas et al. (2022); Lin et al. (2023); Ruffolo et al. (2021). Despite the efforts to measure binding affinity computationally, the reliability and accuracy of these metrics remain an ongoing challenge Johnson et al. (2025).
@@ 2.2 Prior Benchmark Studies Evaluating Binding Affinity via Model Likelihood
Alternatively, some studies retrospectively evaluate machine learning models using historical experimental protein fitness data Chungyoun et al. (2024); Ucar et al. (2024); Wang et al. (2023). These works assume that higher model likelihoods correlate with higher protein fitness, suggesting the model can generate functional sequences. ProteinGym Notin et al. (2023) exemplifies this approach by benchmarking models against experimental data on enzymatic activity, binding, expression, and stability. It measures correlation between model perplexity and fitness scores. However, ProteinGym does not include antibodies. FLAb Chungyoun et al. (2024) addresses this gap by compiling experimental measurements specifically for antibodies, including binding affinity, specificity (e.g., polyreactivity), immunogenicity, and developability metrics. While comprehensive, FLAb does not consider the antigen when evaluating binding affinity. This is a critical limitation, as antibody-antigen interactions are highly specific, and accurate affinity prediction requires modeling the Ab-Ag complex. BindingGYM Lu et al. (2024) is a a related data curation effort for protein-protein interactions. While BindingGYM is a valuable resource, our work differs in scope and construction. We are specifically interested in interactions that involve an antibody’s CDR loops and an antigen and for this reason, only 6 Ab-Ag pairs out of BindgingGYM’s 30 overlap with those we have curated here.
@@ 3.1 Datasets
We compiled 14 datasets containing antigen sequences, antibody heavy and light chain sequences, wild-type complex structures, and experimental binding affinity measurements. (Table 1).
For antibody structure, we focus on variable regions of heavy and light chain, where affinity conferring mutations occur. We only included datasets with at least 20 mutated antibodies to ensure the statistical significance of evaluation.
In each binding affinity study, antibody libraries are constructed through phage or yeast display, introducing mutations via deep mutational scanning or at targeted positions. Some libraries were designed computationally using machine learning Li et al. (2023); Shanehsazzadeh et al. (2023) or biophysical modeling Clark et al. (2006), resulting in variant sets ranging from 67 to 65,535. For consistency, we standardized experimental affinity measurements by taking the negative log for Kd and log for enrichment, so that higher values indicate stronger binding (Supplement LABEL:supplementary:dataset, Table LABEL:tab:dataset_details). While log enrichment is an indirect measure, it reflects how well a variant is retained after antigen-specific selection and correlates with binding strength. When normalized, it provides a scalable proxy for relative binding affinity in high-throughput screens. Notably, AbBiBench is designed to avoid data leakage: while wild-type antibodies and antigens may occur in public datasets such as OAS Olsen et al. (2022a) and SAbDab Dunbar et al. (2014), the specific antibody mutant–antigen complexes curated in our benchmark are not present in any known training corpus.
ID Antibody Antigen Variants Binding score Study
4fqi_h1 CR9114
Influenza A/
New Caledonia/20/99 (H1N1)
HC: 65,094 $\log K_{d}$ Phillips et al. (2021)
4fqi_h3 CR9114
Influenza A/
Wisconsin/67/2005 (H3N2)
HC: 65,535 $\log K_{d}$ Phillips et al. (2021)
3gbn_h1 CR6261
Influenza A/
New Caledonia/20/99 (H1N1)
HC: 1,887 $\log K_{d}$ Phillips et al. (2021)
3gbn_h9 CR6261
Influenza A/
Hong Kong/1073/1999 (H9N2)
HC: 1,842 $\log K_{d}$ Phillips et al. (2021)
aayl49 AAYL49 Spike HR2 HC: 4,312 $\log K_{d}$ Engelhart et al. (2022)
aayl49_ML AAYL49_ML Spike HR2 HC: 8,953 $\log K_{d}$ Li et al. (2023)
aayl50 AAYL50 Spike HR2 LC: 11,473 $\log K_{d}$ Engelhart et al. (2022)
aayl51 AAYL51 Spike HR2 HC: 4,320 $\log K_{d}$ Engelhart et al. (2022)
aayl52 AAYL52 Spike HR2 LC: 13,324 $\log K_{d}$ Engelhart et al. (2022)
2fjg G6.31 VEGF
HC: 2,223
LC: 2,014
log enrichment Koenig et al. (2017)
1mlc D44.1 Hen-egg-white lysozyme
HC: 1,229
LC: 865
log enrichment Warszawski et al. (2020)
1n8z trastuzumab HER2 HC: 419 $\log K_{d}$ Shanehsazzadeh et al. (2023)
1mhp AQC2 Integrin-$\alpha$-1
HC: 37
LC: 25
Both: 5
$\log K_{d}$ Clark et al. (2006)
5a12_ang2 5A12 Ang2
HC: 796
LC: 104
Both: 43
log enrichment Minot & Reddy (2024)
Table 1: Overview of the 14 Ab-Ag binding-affinity assays reported in AbBiBench, showing the number of heavy-chain mutants in each study and the respective binding metric.
@@ 3.2 Protein models
Masked Language Models: Masked protein language models (MLMs) predict masked residues based on context and capture correlations between sequence motifs and higher-level functional properties, enabling their application in antibody design Hie et al. (2024); Meier et al. (2021). We evaluate several representative protein MLMs. ESM-2 Lin et al. (2023) is trained on large-scale protein sequence datasets using a masked language modeling objective. AntiBERTy Ruffolo et al. (2021) is a 12-layer BERT model trained on 57 million heavy- and light-chain sequences from antibody database (OAS Olsen et al. (2022a)). Incorporating structural information into protein language models (PLMs) improves their ability to capture spatial context beyond sequence proximity. SaProt Su et al. (2023) extends ESM-2 with structure-aware tokens from Foldseek Barrio-Hernandez et al. (2023); Van Kempen et al. (2024), embedding residue identity and local structure. ProSST Li et al. (2024) uses geometric vector perceptrons (GVP) encoder Jing et al. (2020) that discretizes local atomic neighborhoods into a compact codebook, with disentangled attention over sequence, structure, and position. ESM-3 Hayes et al. (2025), an upgraded ESM-2, is a multimodal PLM that models sequence, structure, and function through discrete token tracks. Based on such general-purpose protein MLMs, several antibody-specific MLMs have been developed, including CurrAb Burbach & Briney (2025), AbLang Olsen et al. (2022b), and IgBlend Malherbe & Ucar (2024), to capture antibody-specific mutations driven by somatic recombination and hypermutation Ruffolo et al. (2021). CurrAb, a fine-tuned ESM-2, uses curriculum learning to gradually shift from unpaired to paired OAS Olsen et al. (2022a) antibody data while preserving pre-trained knowledge on general proteins.
@@
Autoregressive Protein Language Models: Unlike MLMs, which predict masked tokens based on bidirectional context, autoregressive PLMs generate the next token using only left-to-right context. ProGen2 Nijkamp et al. (2023) is a Transformer‐decoder model that scales up to 6.4 B parameters and is trained on $\sim\!1$ billion natural protein sequences. ProtGPT2 Ferruz et al. (2022) adopts the GPT-2 architecture with 738 M parameters and is trained end-to-end on $\sim\!50$ million protein sequences. Inverse folding models are also autoregressive but based on global structure embeddings. ProGen2 and ProtGPT2 can serve as structure-agnostic comparisons for inverse folding models.
@@
Inverse Folding Models: Inverse folding models aim to predict amino acid sequences from a given protein structure, often in an autoregressive manner. This approach has been applied to antibody mutant design by leveraging known Ab-Ag complex structures Shanker et al. (2024), enabling sequence exploration to identify mutations that preserve or enhance complex stability and binding affinity. Widely used models include ProteinMPNN Dauparas et al. (2022) and ESM-IF Hsu et al. (2022). ProteinMPNN uses a message-passing neural network Gilmer et al. (2017) to model residue interactions before autoregressive sequence generation. ESM-IF1 combines a GVP encoder for extracting backbone-invariant features with a Transformer decoder. AntiFold Høie et al. (2025), based on ESM-IF1, is fine-tuned on antibody structures from SAbDab and OAS.
@@
Diffusion-Based Generative Models: Diffusion models approach antibody design as a denoising process, transforming Gaussian noise into a target antibody by learning the joint distribution of atomic coordinates, orientations, and residue identities. DiffAb Luo et al. (2022) conditions on an antigen–antibody framework complex and jointly diffuses CDR sequence and structure. To assess backbone flexibility, we also evaluated fixed-backbone variants DiffAb_fixbb, a common setting in protein design Anishchenko et al. (2021); Hsu et al. (2022); Ingraham et al. (2019); Luo et al. (2022); Strokach et al. (2020); Tischer et al. (2020). AbDiffuser Martinkus et al. (2023) extends to full-atom generation with physical priors for side chains. IgDiff Cutting et al. (2025) performs de novo backbone generation by sampling variable-region backbones and then fills in the sequences using AbMPNN Dreyer et al. (2023).
@@
CDR Imputation in Geometric Representation: Geometry-aware methods view affinity maturation as filling in missing CDRs within the explicit 3-D Ab-Ag interface. MEAN Kong et al. (2022) masks CDRs on an E(3)-equivariant residue–atom graph containing both chains and epitope; two alternating message-passing blocks jointly restore CDR sequence and backbone. dyMEAN Kong et al. (2023) upgrades this to full-atom, end-to-end design: conserved-framework initialization plus a “shadow paratope” lets the network emit paratope sequence, side-chain geometry, and binding pose in one shot. We also considered fixed-backbone versions of MEAN and dyMEAN, dubbed MEAN_fixbb and dyMEAN_fixbb, respectively.